# Spatial symmetry invariance of solution of Kolmogorov flow

### 23 Mar 2026 v3

@@ Abstract
We prove a mathematical theorem that solution for all $t>0$ of the two-dimensional (2D) Kolmogorov flow governed by Navier-Stokes (NS) equations with periodic boundary condition keeps the same spatial symmetry as its smooth initial condition. This mathematical theorem can be used to check the correctness and reliability of numerical simulations of NS turbulence. For example, it supports the corresponding CNS (clean numerical simulation) results of the 2D turbulent Kolmogorov flow [22, 9] that remain the same spatial symmetry in the whole time interval of simulation, but does not support the corresponding DNS (direct numerical simulation) results that lose the spatial symmetry quickly. In other words, these DNS results violate this mathematical theorem. Thus, this mathematical theorem rigorously confirms that the spatiotemporal trajectories of NS turbulence given by DNS are indeed quickly polluted by numerical noises badly. It also illustrates that CNS can provide helpful enlightenments to deepen our understanding about turbulence and besides approach some mathematical truths about NS equations.
@@ 3 Remarks and discussions
Finally, the theorem of spatial symmetry invariance for 2D Kolmogorov flow can be generalized to the three-dimensional (3D) incompressible Navier-Stokes equations [1], as proved below in the Appendix in a similar way.
@@
Spatial symmetry invariance of solution of the three-dimensional Navier-Stokes equations

### 27 Mar 2026 v4

@@ Abstract
We prove a mathematical theorem that solution for all $t>0$ of the Kolmogorov flow governed by Navier-Stokes (NS) equations with periodic boundary condition keeps the same spatial symmetry as its initial condition. The mathematical proof of a similar theorem for the three-dimensional incompressible NS equations is given in the appendix. These mathematical theorems can be used to check the correctness and reliability of numerical simulations of NS turbulence. For example, they support the corresponding CNS (clean numerical simulation) results of the 2D and 3D turbulent Kolmogorov flows [22, 9, 19] that remain the same spatial symmetry in the whole time interval of simulation, but do not support the corresponding DNS (direct numerical simulation) results that lose the spatial symmetry quickly. In other words, these DNS results violate these mathematical theorems. Thus, these mathematical theorems rigorously confirm that the spatiotemporal trajectories of NS turbulence given by DNS are indeed quickly polluted by numerical noises badly. All of these also indicate that CNS indeed can provide helpful enlightenments to deepen our understanding about turbulence and besides approach some mathematical truths about the NS equations.
@@ 3 Remarks and discussions
Finally, the theorem of spatial symmetry invariance for 2D Kolmogorov flow can be generalized to the three-dimensional (3D) incompressible Navier-Stokes equations [1], as proved below in Appendix in a similar way.
@@
Spatial symmetry invariance of solution of the Navier-Stokes equations